At equilibrium, the concentrations in this system were found to be [N2]=[O2]=0.300 M[N2]=[O2]=0.300 M and [NO]=0.400 M.[NO]=0.400 M.

N2(g)+O2(g)−⇀↽−2NO(g)N2(g)+O2(g)↽−−⇀2NO(g)

If more NONO is added, bringing its concentration to 0.700 M,0.700 M, what will the final concentration of NONO be after equilibrium is re‑established?
Answer #1

Na tę Ž zno [Na] eq = 0.300 : [Q Jea = 0.300 Kc = [NO] =0-400M [NO]? 6.400) cada ] [ ] = 0.300) (0.300) - Kc =1078 No cg) too